# Fine-Grained Cross-View Localization via Local Feature Matching and Monocular Depth Priors

### v1

@@ 5 Conclusion
We propose a simple, accurate, and highly interpretable fine-grained cross-view localization method that matches local features between ground-level and aerial images. By leveraging monocular depth predictors, our method employs scale-aware Procrustes alignment to estimate the ground camera’s location and orientation, and optionally recover the scale between relative depth and metric space. Experiments demonstrate that our approach achieves state-of-the-art performance in challenging scenarios, including cross-area generalization and unknown camera orientation. At inference time, our model can operate with different relative depth predictors without finetuning, with only a minor impact on localization accuracy. The combination of this flexibility and consistently strong localization performance makes our method well-suited for real-world deployment.
In this Appendix, we provide the following information to support the main paper:
@@
  2. B.
    Additional results on using relative depth in inference
  3. C.
@@
  7. G.
    Limitations.
  8. H.
    Broader impact.
Appendix A Additional Local Feature Matching Results
@@
Success cases: In Fig. 5 (a)–(f), we demonstrate that our model effectively matches various types of road markings across views, including painted text (a)–(d) and lines (e)–(f). Notably, in Fig. 5 (b), the ‘bus’ marking is not clearly visible in the aerial view, making it difficult for humans to identify. Yet, our model successfully matches it. In general, road markings are distinctive across views, and accurately matching them significantly aids localization. When road markings are repetitive, e.g., on highways (Fig. 5 (h)), our method leverages other landmarks, such as overhead road signs and streetlights, to help anchor the keypoint locations. In urban areas, there is typically rich visual information available for matching across ground and aerial views (Fig. 5 (i)–(l)), including buildings, zebra crossings, and lane lines. Our model effectively exploits all of these cues for localization.
@@
Failure cases: Typical failure cases are shown in Fig. 6. When there is dense vegetation (Fig. 6 (a)–(c)), the matchable information across views is very limited, making precise localization extremely challenging, even for humans. Interestingly, in Fig. 6 (b), our method still correctly matches the gap between two buildings in the ground view to the corresponding gap in the aerial view. In Fig. 6 (c), our method predicts a more reasonable position aligned with the path seen in the ground image, whereas the ground-truth location appears incorrect, placing the camera on the grass rather than on the path. In Fig. 6 (d), the two narrow roads with trees and a zebra crossing in front appear visually similar. At the true location, a tree partially blocks the zebra crossing in the aerial view, making it harder to distinguish. As a result, our method matches the zebra crossing to an incorrect one, leading to an erroneous location estimate. In Fig. 6 (e), although our method correctly matches the pole between the ground and aerial views, most of the scene is featureless and contains many repetitive patterns, making localization difficult. Similarly, in Fig. 6 (f), it is nearly impossible to determine the exact location over the water due to the lack of distinctive features.
@@
Refer to caption(e)
Refer to caption(f)
@@
Appendix B Additional Results on Using Relative Depth in Inference
We conduct additional experiments on using relative depth at inference time for our model (trained with metric depth predictions from UniK3D Piccinelli et al. (2025)). Specifically, we use BiFuse++ Wang et al. (2022) as our relative depth predictor for these experiments.
Initial scale and maximum depth threshold: As mentioned in Sec. 4.2 of the main paper, when using relative depth, we apply an initial scaling factor to all relative depth maps to bring the values into a reasonable depth range, determined by visually inspecting a few examples. We then apply a predefined maximum depth threshold, the same one used for the metric depth model, to filter out matches corresponding to sky and distant objects. In practice, the initial scale and maximum depth threshold can be determined for each dataset by visually inspecting a few examples. Here, we study how consistent the model’s predictions are when these settings are varied.
First, we evaluate the model’s robustness against different scales of the relative depth. We apply an additional scaling factor on top of the initial scale. To ensure the same objects are kept for pose estimation, the same scaling factor is also applied to the maximum depth threshold. As shown in Fig. 7, the test errors remain stable across different scaling factors during inference. This observation is consistent with our findings on applying scaling factors to metric depth in the main paper (Sec. 4.3). Therefore, in practice, when using relative depth, people do not need to carefully select an initial scale.
@@ Appendix B Additional Results on Using Relative Depth in Inference
Figure 7: VIGOR test errors when different scaling factors are applied to the relative depth during inference. We use the model trained with metric depth from UniK3D Piccinelli et al. (2025), with unknown orientation. At inference time, we use relative depth from BiFuse++ Wang et al. (2022).
Table 4: Robustness to different maximum depth thresholds. We apply different maximum depth thresholds during inference to remove sky and distant objects (after applying the initial scale to the relative depth from BiFuse++ Wang et al. (2022)). Best results are shown in bold. During training, we use metric depth predictions from UniK3DPiccinelli et al. (2025), with a maximum depth threshold of 35 m.
@@
Scale recovery: Next, we present more results on scale recovery. Here, we apply an arbitrary initial scale to the relative depth and visualize the bird’s-eye view (BEV) points both before and after applying our estimated scale. Fig. 8 and 9 illustrate success cases. Before applying the estimated scale (third row), the BEV points clearly do not align with the aerial image. After applying the estimated scale (fourth row), the BEV points align well with the contents in the aerial image, such as building facades and road markings. We observe that scale recovery is often reliable when the estimated pose is accurate, as both are directly derived from local feature matching, and an accurate pose results from precise local feature correspondences.
Next, we show the failure cases in Fig. 10. Sometimes, the misalignment between BEV points and the aerial image is due to the error in the estimated pose. For example, the first column of Fig. 10 shows that the scene has a symmetric layout, and our method outputs an incorrect pose by placing the BEV points at a symmetric but incorrect location. Interestingly, the transformed BEV points still align well with the aerial images, such as the zebra crossing. In the last column, the ground truth position lies on the zebra crossing, whereas the ground image indicates there is still some distance to it. Although our estimated pose deviates from the ground truth, the transformed BEV points appear to align well with the aerial image, suggesting a plausible localization despite the numerical error.
Figure 8: Success cases of scale recovery. BEV points overlaid on the aerial image after applying the predicted rotation, translation, and either the initial or predicted scale transformations. Third row: using an arbitrary initial scale. Fourth row: using the predicted scale.
Figure 9: Success cases of scale recovery. BEV points overlaid on the aerial image after applying the predicted rotation, translation, and either the initial or predicted scale transformations. Third row: using an arbitrary initial scale. Fourth row: using the predicted scale.
Figure 10: Failure cases of scale recovery. BEV points overlaid on the aerial image after applying the predicted rotation, translation, and either the initial or predicted scale transformations. Third row: using an arbitrary initial scale. Fourth row: using the predicted scale.
Appendix C Training and Testing with Relative Depth
As mentioned in Sec. 4.1 of the main paper, we also experimented with training and testing using only relative depth. This setup is more challenging, since inducing correct correspondences on the BEV plane requires both ground-truth pose and scale (i.e., metric depth). As a result, the infoNCE loss cannot be directly applied and we lack direct supervision of the correspondences during training. We then explored two settings. (1) Training without infoNCE losses, i.e., setting the weight $\beta=0$. (2) With pseudo ground truth for infoNCE losses: At each training iteration, we estimate the scale and use it to generate pseudo ground-truth positive matches for the infoNCE losses.
@@ Appendix C Training and Testing with Relative Depth
Table 5: VIGOR test results. Best in bold. (1) and (2): Training and testing with relative depth predicted by BiFuse++Wang et al. (2022). For reference, we also include the model trained and tested with metric depth from Unik3DPiccinelli et al. (2025), as well as the model trained with metric depth and tested with relative depth.
@@
Appendix D Additional Proof for Scale-Aware Procrustes Alignment
As discussed in Sec. 3.2 of the main paper, we estimate pose and scale using scale-aware Procrustes alignment Umeyama (1991). Here, we provide the proof of the two main steps involved:
@@ Appendix D Additional Proof for Scale-Aware Procrustes Alignment
    The rotation from ground points to aerial points is computed as $\mathbf{R}=\mathbf{V}\mathbf{U}^{\top}$.
@@
    When the ground points lie in the same metric space as the aerial points, i.e., there is no scale difference between them, we have $\operatorname{Tr}(\mathbf{\Sigma})=\sum_{n=1}^{N}w_{n}\left\|\tilde{\mathbf{P}}_{n}\right\|^{2}$.
First, we recall the key notations: The inputs to the scale-aware Procrustes alignment are the coordinates of ground–aerial correspondences along with their weights. We denote the ground points as $\mathbf{P}$, the aerial points as $\mathbf{Q}$, and the weight of the $n$-th correspondence as $w_{n}$. The objective is to find the optimal rotation $\mathbf{R}$, translation $\mathbf{t}$, and scale $s$ such that $\mathbf{Q}=s(\mathbf{R}\cdot\mathbf{P})+\mathbf{t}$. We denote the weighted centroids of the aerial and ground point sets as $\bar{\mathbf{Q}}$ and $\bar{\mathbf{P}}$, and the centered point sets as $\tilde{\mathbf{Q}}$ and $\tilde{\mathbf{P}}$, respectively.
@@
In Eq. 7, the term $\left\|\mathbf{R}\,\cdot\tilde{\mathbf{P}}_{n}-\tilde{\mathbf{Q}}_{n}\right\|^{2}$ can be rewritten as (for simplicity, we omit the explicit dot product notation):
@@
$=\left\|\tilde{\mathbf{P}}_{n}\right\|^{2}+\left\|\tilde{\mathbf{Q}}_{n}\right\|^{2}-2(\tilde{\mathbf{Q}}_{n})^{\top}\mathbf{R}\tilde{\mathbf{P}}_{n}.$
Note that $\left\|\tilde{\mathbf{P}}_{n}\right\|^{2}$ and $\left\|\tilde{\mathbf{Q}}_{n}\right\|^{2}$ are fixed, and finding the optimal rotation effectively depends only on the term $(\tilde{\mathbf{Q}}_{n})^{\top}\mathbf{R}\tilde{\mathbf{P}}_{n}$. Then, the objective in Eq. 7, which sums over each point correspondence, can be rewritten in matrix form as:
@@
where $\mathbf{W}$ is a diagonal matrix with each element $\mathbf{W}_{n,n}=w_{n}$. In Eq. 10, the matrix product $\tilde{\mathbf{P}}\mathbf{W}\tilde{\mathbf{Q}}^{\top}$ constructs a $2\times 2$ square matrix $\mathbf{C}$, which can be decomposed using singular value decomposition (SVD) as $\mathbf{C}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$. Accordingly, the objective function in Eq. 10 can be reformulated as:
@@
The centered target points $\tilde{\mathbf{Q}}_{n}$ can be obtained by rotating and scaling the centered source points $\tilde{\mathbf{P}}_{n}$, satisfying $\tilde{\mathbf{Q}}_{n}=s\mathbf{R}\tilde{\mathbf{P}}_{n}$. Since we have already estimated the rotation $\mathbf{R}=\mathbf{V}\mathbf{U}^{\top}$, Eq. 13 can be formulated as
@@
Appendix E Additional Details on InfoNCE Losses

### v2

@@ 5 Conclusion
We propose a simple, accurate, and interpretable fine-grained cross-view localization method that matches local features between ground and aerial images. Our method learns correspondences from camera pose supervision and leverages monocular depth predictors with scale-aware Procrustes alignment to estimate the camera pose and recover the scale of relative depth. Experiments demonstrate state-of-the-art performance in challenging scenarios, including cross-area generalization and unknown camera orientation. Furthermore, our local feature matching enhances interpretability, enabling visual identification of erroneous predictions as well as RANSAC-based outlier detection.
Appendix
Here we provide supplementary material to support the main paper:
@@
    Additional ablation study.
  2. B.
    Test on non-road facing samples on KITTI
  3. C.
@@
  4. D.
    Additional results on using relative depth in inference.
  5. E.
@@
Appendix A Additional ablation study
We conduct an additional ablation study on the VIGOR same-area validation set with unknown orientation, analyzing the effects of the number of sampled correspondences $N$, the number of aerial points, the aerial image resolution, and the use of RANSAC.
Number of sampled correspondences $N$: Our default setting uses $N=1024$ (following Xia & Alahi (2025)) during training. In inference, if RANSAC is used, we sample only 3 correspondences per RANSAC loop to compute the pose and scale. When RANSAC is not used, we use the same $N=1024$.
We evaluate two cases: 1) varying $N$ only in inference: a model trained with $N=1024$ is evaluated with $N=256,512,1024,$ and $2048$ (without RANSAC). 2) varying $N$ in both training and inference: validation results (with RANSAC) of models trained with $N=256,512,1024,$ and $2048$.
As shown in Tab. 4, the number of correspondences $N$ does not have a strong impact on performance. When varying $N$ during inference, the difference in mean localization error is less than 0.2 m for $N=512,1024,2048$. The error increases more noticeably when a smaller $N$, such as 256, is used. Similarly, the model remains robust to different $N$ during training, as long as $N$ is not too small (e.g., $N\leq 256$).
Table 4: Ablation on the number of sampled correspondences NN. Best in bold..
Mode N Localization (m) Orientation ()
Mean ↓ Median ↓ Mean ↓ Median ↓
Inference
2048 4.66 2.88 11.13 3.92
1024 (default) 4.60 2.81 11.19 4.04
512 4.71 2.91 12.06 4.20
256 4.90 2.98 12.97 4.78
Training
2048 3.91 1.76 9.25 1.73
1024 (default) 3.86 1.75 9.30 1.93
512 4.12 1.80 9.71 1.80
256 4.31 1.90 11.27 2.17
Number of aerial points: Next, we examine how the number of aerial points affects performance. Our default setting, following Xia & Alahi (2025), uses $41\times 41$ aerial points for fair comparison. As shown in Tab. 5, our model is robust to small variations in the number of aerial points. When more than $36\times 36$ points are used, the change in mean localization error is less than 0.2 m. However, when the number of points is significantly reduced (e.g., $31\times 31$), performance degrades noticeably.
Table 5: Ablation on the number of aerial points. Best in bold.
Number of Aerial Points Localization (m) Orientation ()
Mean ↓ Median ↓ Mean ↓ Median ↓
$31{\times}31$ 4.27 2.16 10.01 2.49
$36{\times}36$ 3.98 1.88 9.38 2.10
$41{\times}41$ (default) 3.86 1.75 9.30 1.93
$46{\times}46$ 3.86 1.72 9.31 1.91
$51{\times}51$ 3.82 1.70 9.18 1.90
$56{\times}56$ 3.91 1.68 9.61 1.97
Aerial image resolution: We also study the effect of varying aerial image resolution in inference. As shown in Tab. 6, our model is robust to such variations. All aerial images are first processed by the pre-trained DINOv2 feature extractor. Regardless of input resolution, we consistently sample a fixed $41\times 41$ grid of aerial points from the feature map. Since a downsampled aerial image still covers the same geographic area, the sampled points correspond to the same geo-locations, independent of resolution. Thus, any effect on pose estimation accuracy mainly stems from changes in DINOv2 features due to downsampling in lower-resolution inputs.
Table 6: Ablation on aerial image resolution. Best in bold.
Resolution Localization (m) Orientation ()
Mean ↓ Median ↓ Mean ↓ Median ↓
$630{\times}630$ (default) 3.86 1.75 9.30 1.93
$574{\times}574$ 3.77 1.71 9.52 1.95
$518{\times}518$ 3.82 1.76 9.17 1.94
$448{\times}448$ 3.90 1.77 9.49 1.97
$392{\times}392$ 3.99 1.84 10.22 2.13
$336{\times}336$ 4.17 1.96 10.36 2.28
RANSAC: Finally, we report results of in inference with or without RANSAC . As shown in Tab. 7, using RANSAC yields a $\sim$0.74 m reduction in mean localization error.
Table 7: Ablation on RANSAC. Best in bold.
RANSAC Localization (m) Orientation ()
Mean ↓ Median ↓ Mean ↓ Median ↓
Without RANSAC 4.60 2.81 11.19 4.04
With RANSAC 3.86 1.75 9.30 1.93
Appendix B Test on non-road facing samples on KITTI
In Sec. 4.3, we noted that our method computes orientation from correspondences, making it difficult to exploit the prior that most ground images in KITTI are aligned with the road direction. Here, we provide an additional quantitative comparison between our method and CCVPE on manually selected KITTI samples that are not aligned with the road direction, denoted as non-road.
As shown in Tab. 8, our method consistently outperforms CCVPE (which directly regresses orientation and can easily capture the road-facing bias). This setting better reflects real-world scenarios, where the orientation can be arbitrary and no pre-processing (such as rotating the aerial image for rough alignment) is required. When comparing the “non-road” subset to the full test set (Tab. 1), both CCVPE and our method exhibit increased orientation errors. However, the increase is significantly smaller for our method, indicating greater robustness to challenging viewpoints. Interestingly, we observe that localization errors in the non-road subset are lower than those on the full set. This is likely due to reduced location ambiguity at intersections compared to the more monotonous structure of road-aligned scenes.
Table 8: Test on non-road facing samples on KITTI. Best in bold.
Orientation Setting Localization (m) Orientation ()
Test Set Method Mean ↓ Median ↓ Mean ↓ Median ↓
Unknown
Cross-area (non-road) CCVPE 8.33 5.99 89.78 93.97
Cross-area (non-road) Ours 7.04 5.00 58.97 47.98
Same-area (non-road) CCVPE 3.19 1.81 26.10 16.88
Same-area (non-road) Ours 1.17 1.01 9.79 8.00
Appendix C Additional Local Feature Matching Results
@@
Success cases: Fig. 6 presents results on VIGOR. In (a)–(f), we demonstrate that our model effectively matches various types of road markings across views, including painted text (a)–(d) and lines (e)–(f). In general, road markings are distinctive across views, and accurately matching them significantly aids localization. Besides road markings, our method also leverages other landmarks, such as overhead road signs and streetlights, to help anchor the keypoint locations.(Fig. 6 (g)–(h)). In urban areas, there is typically rich visual information available for matching across ground and aerial views (Fig. 6 (i)–(l)), including buildings, zebra crossings, and lane lines. Our model effectively exploits all of these cues for localization.
Fig. 7 presents results on KITTI. Our method primarily leverages structures such as roads, sidewalks, and fences in front of the vehicle to establish correspondences with the aerial view.
@@
Failure cases: Typical failure cases are shown in Fig. 8. When there is dense vegetation (Fig. 8 (a),(b)), the matchable information across views is very limited, making precise localization extremely challenging, even for humans. Interestingly, in Fig. 8 (b), our method still correctly matches the gap between two buildings in the ground view to the corresponding gap in the aerial view. In Fig. 8 (c), the two narrow roads with trees and a zebra crossing in front appear visually similar. At the true location, a tree partially blocks the zebra crossing in the aerial view, making it harder to distinguish. As a result, our method matches the zebra crossing to an incorrect one, leading to an erroneous location estimate. Finally, in Fig. 8 (f), it is nearly impossible to determine the exact location over the water due to the lack of distinctive features.
@@
Figure 7: Success cases for localization on the KITTI same-area test set under $\pm 10^{\circ}$ orientation noise. We visualize the top 50 correspondences, ranked by matching score
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
@@
Appendix D Additional Results on Using Relative Depth in Inference
We conduct additional experiments on using relative depth at inference time for our model (trained with metric depth predictions from UniK3D (Piccinelli et al., 2025)). Specifically, we use BiFuse++ (Wang et al., 2022) as our relative depth predictor for these experiments.
Initial scale and maximum depth threshold: As mentioned in Sec. 4.2, when using relative depth, we apply an initial scaling factor to all relative depth maps to bring the values into a reasonable depth range, determined by visually inspecting a few examples. We then apply a predefined maximum depth threshold, the same one used for the metric depth model, to filter out matches corresponding to sky and distant objects. In practice, the initial scale and maximum depth threshold can be determined for each dataset by visually inspecting a few examples. Here, we study how consistent the model’s predictions are when these settings are varied.
First, we evaluate the model’s robustness against different scales of the relative depth. We apply an additional scaling factor on top of the initial scale. To ensure the same objects are kept for pose estimation, the same scaling factor is also applied to the maximum depth threshold. As shown in Fig. 9, the test errors remain stable across different scaling factors during inference. This observation is consistent with our findings on applying scaling factors to metric depth in Sec. 4.3. Therefore, in practice, when using relative depth, people do not need to carefully select an initial scale.
@@ Appendix D Additional Results on Using Relative Depth in Inference
Figure 9: VIGOR test errors when different scaling factors are applied to the relative depth during inference. We use the model trained with metric depth from UniK3D (Piccinelli et al., 2025), with unknown orientation. At inference time, we use relative depth from BiFuse++ (Wang et al., 2022).
Table 9: Robustness to different maximum depth thresholds. We apply different maximum depth thresholds during inference to remove sky and distant objects (after applying the initial scale to the relative depth from BiFuse++ (Wang et al., 2022)). Best results are shown in bold. During training, we use metric depth predictions from UniK3D(Piccinelli et al., 2025), with a maximum depth threshold of 35 m.
@@
Ground-aerial layout alignment: Next, we present additional results on ground–aerial layout alignment. We visualize the bird’s-eye view ground layout after applying our estimated pose and scale. As shown in Fig. 10, the quality of alignment directly reflects localization accuracy: when the alignment is precise, the estimated pose is accurate; when the alignment is off, as in the last example, localization also fails. This interpretability provides a practical means of identifying potentially erroneous predictions.
Figure 10: Ground layout overlaid on the aerial image after applying the predicted rotation, translation, and scale transformations. The alignment directly reflects localization quality: the first three examples show successful localization, while the last one illustrates a failure case. Relative depth is obtained from BiFuse++ (Wang et al., 2022).
Appendix E Training and Testing with Relative Depth
As mentioned in Sec. 4.3, we also experimented with training and testing using only relative depth. This setup is more challenging, since inducing correct correspondences on the BEV plane requires both ground-truth pose and scale (i.e., metric depth). As a result, the infoNCE loss cannot be directly applied and we lack direct supervision of the correspondences during training. We then explored two settings. (1) Training without infoNCE losses, i.e., setting the weight $\beta=0$. (2) With pseudo ground truth for infoNCE losses: At each training iteration, we estimate the scale and use it to generate pseudo ground-truth positive matches for the infoNCE losses.
@@ Appendix E Training and Testing with Relative Depth
Table 10: VIGOR test results. Best in bold. (1) and (2): Training and testing with relative depth predicted by BiFuse++(Wang et al., 2022). For reference, we also include the model trained and tested with metric depth from Unik3D(Piccinelli et al., 2025), as well as the model trained with metric depth and tested with relative depth.
@@
Appendix F Additional Proof for Scale-Aware Procrustes Alignment
As discussed in Sec. 3.2 of the main paper, we estimate pose and scale using scale-aware Procrustes alignment (Umeyama, 1991) from local correspondences. Let $\{(\mathbf{P}_{n},\mathbf{Q}_{n},w_{n})\}^{N}_{n=1}$ denote weighted correspondences between ground points $\mathbf{P}$ and aerial points $\mathbf{Q}$ with weights $w_{n}\geq 0$ and $\sum_{n}w_{n}=1$. We denote the weighted centroids as $\bar{\mathbf{P}}$ and $\bar{\mathbf{Q}}$, and the centered point sets as $\tilde{\mathbf{P}}=\mathbf{P}-\bar{\mathbf{P}}$ and $\tilde{\mathbf{Q}}=\mathbf{Q}-\bar{\mathbf{Q}}$. The objective is to find the optimal rotation $\mathbf{R}$, translation $\mathbf{t}$, and scale $s$ satisfying,
$\mathbf{Q}=s(\mathbf{R}\cdot\mathbf{P})+\mathbf{t}.$ (7)
Below we provide proofs for two key steps:
@@ Appendix F Additional Proof for Scale-Aware Procrustes Alignment
    The optimal rotation from ground points to aerial points is $\mathbf{R}=\mathbf{V}\mathbf{U}^{\top}$.
@@
    When ground and aerial points lie in the same metric space, i.e., there is no scale difference between them, we obtain $\operatorname{Tr}(\mathbf{\Sigma})=\sum_{n=1}^{N}w_{n}\left\|\tilde{\mathbf{P}}_{n}\right\|^{2}$.
@@
The term $\left\|\mathbf{R}\,\cdot\tilde{\mathbf{P}}_{n}-\tilde{\mathbf{Q}}_{n}\right\|^{2}$ can be rewritten as (for simplicity, we omit the explicit dot product):
@@
$=\left\|\tilde{\mathbf{P}}_{n}\right\|^{2}+\left\|\tilde{\mathbf{Q}}_{n}\right\|^{2}-2\tilde{\mathbf{Q}}_{n}^{\top}\mathbf{R}\tilde{\mathbf{P}}_{n}.$
Note that $\left\|\tilde{\mathbf{P}}_{n}\right\|^{2}$ and $\left\|\tilde{\mathbf{Q}}_{n}\right\|^{2}$ are fixed, so the optimal rotation depends only on the inner product $\tilde{\mathbf{Q}}_{n}^{\top}\mathbf{R}\tilde{\mathbf{P}}_{n}$. Hence, the objective in Eq. 8, which sums over all correspondences, can be written in matrix form as:
@@
where $\mathbf{W}$ is a diagonal matrix with entrie $\mathbf{W}_{n,n}=w_{n}$. In Eq. 11, the matrix product $\tilde{\mathbf{P}}\mathbf{W}\tilde{\mathbf{Q}}^{\top}$ constructs a $2\times 2$ square matrix $\mathbf{C}$, which can be decomposed using singular value decomposition (SVD) as $\mathbf{C}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$. Accordingly, the objective function in Eq. 11 can be reformulated as:
@@
The centered target points $\tilde{\mathbf{Q}}_{n}$ can be obtained by rotating and scaling the centered source points $\tilde{\mathbf{P}}_{n}$, satisfying $\tilde{\mathbf{Q}}_{n}=s\mathbf{R}\tilde{\mathbf{P}}_{n}$. With estimated rotation $\mathbf{R}=\mathbf{V}\mathbf{U}^{\top}$, Eq. 14 can be formulated as
@@
Appendix G Additional Details on InfoNCE Losses